# Multimode Entangled Squeezed Light Generation and Propagation in a Coupled-Cavity Photonic Crystal

### v2

@@ III.1 Coupled-Cavity System QMs
As our first step, we calculate the QMs of the individual defects and present the tight-binding formalism that we use to obtain the leaky QMs of the entire PC CCS discussed earlier and shown in Fig. 1. We first discuss some general properties of PC slabs and defect modes, and then show how we determine the QMs of a general system of coupled defects. We then compute the QMs of the full system and each of the subsystems.
III.1.1 General Procedure
@@
Defect-free PC slabs are periodic dielectric structures which, if properly designed, can support a photonic band gap - a set of frequencies for which light cannot propagate in the PC - similar to an electronic band gap in which no electronic states exist [32]. Both of these band gaps arise from the periodicity of the underlying structure. The photonic band structure (and thus the size and location of the photonic band gap) of a PC slab is easily calculated using finite-difference time-domain (FDTD) simulations or finite-element analysis (FEA) by exploiting the high symmetry in a defect-free PC slab.
@@
One may form single-defect cavities in a PC slab by removing or shrinking an air hole at a particular lattice site, and thus locally disrupting the periodicity of the structure. Defects of this type are displayed in Fig. 1. These single-defect cavities in a square-lattice PC slab support a quadrupole-like resonant QM localized to the lattice site of the altered air hole [29]. The resonant frequency of the single-defect cavity modes formed in this manner lies within the transverse electric photonic band gap of the underlying PC slab.
@@
The positive frequency part of the electric field of light in the $q^{\text{th}}$ individual defect QM is given by [30]
@@
××$\curl\curl\textbf{M}_{q}(\textbf{r})-\frac{\tilde{\Omega}_{q}^{2}}{c^{2}}\epsilon_{q}(\textbf{r})\textbf{M}_{q}(\textbf{r})=0,$ (3)
@@
Figure 4: Single-defect QM frequency and quality factor as a function of defect air hole radius for the chosen PC. Computed with ANSYS Lumerical FDTD [33].
The $\textbf{M}_{q}(\textbf{r})$ act as a basis for full system QMs that have frequencies well inside the chosen band gap of the PC. We use a tight-binding approach and expand the system QMs, $\textbf{N}_{\mu}(\textbf{r})$, in terms these as [27]
@@
××$\curl\curl\textbf{N}_{\mu}(\textbf{r})-\frac{\tilde{\omega}_{\mu}^{2}}{c^{2}}\epsilon(\textbf{r})\textbf{N}_{\mu}(\textbf{r})=0,$ (6)
@@
Eqs. (3), (4), and (6) lead to the generalized eigenvalue problem [27]
@@
$\bra{\textbf{N}_{\mu}}\ket{\textbf{N}_{\nu}}$ $\equiv\int_{\infty}d^{3}\textbf{r}\epsilon(\textbf{r})\textbf{N}_{\mu}^{*}(\textbf{r})\cdot\textbf{N}_{\nu}(\textbf{r})$ (10)
@@
where $V_{\mu}$ is the $\mu^{\text{th}}$ column of $\stackrel{{\scriptstyle\leftrightarrow}}{{V}}$. The $O_{\mu\nu}$ are elements of the square Hermitian matrix O, called the QM overlap matrix. Note that although they are chosen to be normalized ($O_{\mu\mu}=1\ \forall\ \mu$), the full system QMs are not orthogonal. This is due to the loss in the system, which makes the problem non-Hermitian [29, 30]. In Fig. 5 we plot the overlaps between all of the 184 full-system QMs, as computed with Eq. (12). It is immediately apparent that a significant number of QM pairs are far from being orthogonal. In particular, some of the QMs that are very close in frequency have overlaps that are significantly more than 0.1.
The non-orthogonality of the full-system QMs needs to be accounted for in the commutation relations of the system QM creation and annihilation operators. As we will show in Sec. III, the effect of the overlap matrix mathematically is to mix the QMs during the free evolution of the light. This mixing will have non-trivial effects on the dynamics of the system and cannot be ignored. We will address the consequences of the mon-orthogonality in detail in Sec. IV. We now proceed with the computation of the QMs in each of the subsystems.
III.1.2 Resonant Structure Quasi-Modes
@@
××$\curl\curl\textbf{C}_{j}(\textbf{r})-\frac{\tilde{\omega}_{j}^{2}}{c^{2}}\epsilon_{\text{RS}}(\textbf{r)}\textbf{C}_{j}(\textbf{r})=0,$ (14)
@@
In this case the RS is constructed from three single-defect cavities arranged in a line (see Fig. 1). Following Section III.1.1, the tight-binding eigenvalue problem for the RS QMs is
@@
In Fig. 6, we plot $|C_{j,x}(x,y,0)|^{2}$ for each of the three RS QMs and list their complex frequencies. We can see that the three RS QMs strongly overlap in space. The real parts of the frequencies of the RS QMs dictate the operating frequencies for SFWM. The band structures of the CROWs plotted in Fig. 2 are designed to center on the RS QM frequencies. The hole sizes of the defects used to arrive at the final frequency configuration were chosen by comparing the computed QMs of combinations of three defect configurations in which the wing cavities were identical.
[…]
Figure 6: Plot of $|C_{j,x}(x,y,0)|^{2}$ for the three RS QMs. The area of each subfigure is enclosed by the purple rectangle in Fig. 1. The color scale is linear, with bright yellow indicating the peak value and dark blue indicating zero. The complex frequency for each of the QMs (when the CROWs are absent) are given below each plot.
III.1.3 Bloch Modes of an Infinite CROW
The remaining subsystems are the three CROWs. To understand the modes of the finite CROWS in the system, it is useful to first determine the Bloch QMs of infinite CROWs. In the nearest-neighbor tight-binding (NNTB) approximation, these modes can be expanded in the basis of single-defect cavity QMs as [27]
@@ III.1.3 Bloch Modes of an Infinite CROW
where $d=3a$ is the CROW period and $N$ is the number of cavities in the CROW, which is taken to be of length $L=Nd$. In the limit that $N\rightarrow\infty$, the solution to the generalized eigenvalue problem for the CROW in the NNTB gives [25]
@@
III.2 Pump Evolution
Now that we have computed the various bases with which we will describe the light, we turn our attention to modeling the evolution of the pump light. The pump at time $t=0$ starts in a coherent (classical) state in the pump CROW. We choose the initial state of the pump to be a Gaussian in the Bloch modes, centered at $k_{0}=-\frac{\pi}{2d}$. With this choice, the propagation speed is maximized, there is no waveguide group velocity dispersion, and the frequency is centered at the pump frequency of the RS: $\omega_{k_{0}}=\Omega_{0}=\omega_{P}$. Thus, the positive frequency part of the electric field of the pump light at time $t=0$ is given by
@@ III.2 Pump Evolution
Now, to determine the time evolution of the pump in the full system, we need to expand the pump field at $t=0$ in terms of the QMs of the full system, since we know that these evolve in time according to Eq. (5). We start by inverting Eq. (4) to obtain the single cavity modes in terms of the modes of the full system:
@@
III.3 Nonlinear Generation
In this section, we present our approach to calculating the density operator of the two-mode squeezed thermal state generated in the RS of our system via SFWM. We make the undepleted pump approximation, i.e., we assume that the pump is unaffected by the nonlinear interaction. This is a good approximation as long as the number of photons generated is much less than the number of photons in the pump in the RS, which (as we shall show) is the case in our simulations. We shall label the ladder operators for the $q^{th}$ single-defect cavity mode as $\quantity{a_{q},\ a^{\dagger}_{q}}$. Restricting ourselves to the three modes of the RS, we define the ladder operators, $\{d_{j},\ d_{j}^{\dagger}\}$, for the $j^{th}$ RS QM in terms of the single-defect QM ladder operators as:
$d_{j}=\sum_{q=1}^{3}u_{qj}a_{q},\quad d_{j}^{\dagger}=\sum_{q=1}^{3}u^{*}_{qj}a_{q}^{\dagger}.$ (27)
In this work, we only consider the SFWM nonlinear interaction in which two pump photons are exchanged for one signal and one idler photon. The Hamiltonian for the SFWM process is given by [11, 23]
$H^{\text{RS}}=H^{\text{RS}}_{\text{L}}+H^{\text{RS}}_{\text{NL}},$ (28)
where
is the linear part of the Hamiltonian and
is the nonlinear part [34, 12], where $\alpha_{P}(t)$ is the dimensionless time-dependent coherent state amplitude of the pump in the pump mode of the RS,
×$\times\sum_{i,j,k,l=1}^{3}\int d^{3}\textbf{r}\chi^{(3)}_{ijkl}(\textbf{r})\epsilon_{\text{RS}}(\textbf{r})$
×$\times C_{P,i}(\textbf{r})C_{P,j}(\textbf{r})C^{*}_{S,k}(\textbf{r})C^{*}_{I,l}(\textbf{r}),$ (31)
is the effective nonlinear parameter carrying the dimensions of a rate, and $\chi^{(3)}_{ijkl}$ are the elements of the third-order nonlinear susceptibility tensor for the dielectric substrate, which is assumed to be homogeneous in the substrate and zero elsewhere. Using $\chi_{iiii}^{(3)}=4n_{2}n^{2}\epsilon_{0}c/3$ with $n_{2}=4.5\times 10^{-18}\ \text{m}^{2}\cdot\text{W}^{-1}$ for silicon at 1550 nm [35, 36], and setting all other components to zero, we find that $\chi_{\text{eff}}\approx(1.84+0.07i)\times 10^{5}\ \text{s}^{-1}$.
The time-dependent pump amplitude, $\alpha_{P}(t)$, is obtained by projecting $\textbf{E}_{P}^{+}(\textbf{r},t)$ onto the pump QM of the RS  [^1] The second line is approximate since in the previous line we use $\epsilon_{\text{RS}}$, while $\stackrel{{\scriptstyle\leftrightarrow}}{{B}}$ was computed using $\epsilon$.:
@@ III.3 Nonlinear Generation
The dynamics of the generated light in the RS described by the Hamiltonian of Eq. (28) is obtained by solving the Lindblad master equation for the density operator projected onto the RS QMs, $\rho^{\text{RS}}$ [28]. The Lindblad master equation takes the form [28]
[…]
where the $\Gamma_{j}$ are the loaded power decay rates of the signal $(j=S)$ and idler $(j=I)$ modes (i.e., the decay rates when the RS is coupled to all three CROWs), and we assume that the RS QMs are orthogonal, which is an excellent approximation for our RS ($|O_{SI}|\leq 0.01)$.
Vendromin and Dignam have shown that the solution to Eq. (33) is the density operator of a two-mode STS [12], which takes the form
$\rho^{RS}=S_{2}(\xi)\rho_{th}(t)S_{2}^{\dagger}(\xi),$ (34)
where
$S_{2}(\xi)=\exp(\xi^{*}d_{S}d_{I}-\xi d_{S}^{\dagger}d_{I}^{\dagger})$ (36)
is the two-mode squeezing operator with time-dependent squeezing parameter $\xi(t)=r(t)e^{i\theta(t)}$, where $r(t)$ is the squeezing amplitude and $\theta(t)$ is the squeezing phase.
Writing the nonlinear driving term as
@@
we find, for our choice of initial pulse in the structure we have designed, that $\beta$ is independent of $t$ over the time range for which the pump is non-negligible. When $\beta$ is independent of $t$ and the initial state is the vacuum state, it can be shown that the squeezing phase evolves simply according to [12]
@@
and the squeezing amplitude and thermal photon numbers are governed by the following set of coupled first-order differential equations [12]:
@@
where we employ the dimensionless time units $\tilde{t}=\Gamma_{+}t$ with $\Gamma{\pm}=(\Gamma_{S}\pm\Gamma_{I})/2$. The parameter $\zeta=\Gamma_{-}/\Gamma_{+}$, quantifies the difference in the loss rates in the two modes. Finally,
$g(\tilde{t})=\frac{2\absolutevalue{\alpha_{P}^{2}(\tilde{t})\chi_{\text{eff}}}}{\Gamma_{+}}$ (42)
is the pump strength in the pump mode. It is essentially the ratio of intensity buildup of the light in the pump mode to the decay rate of generated light, $\Gamma_{+}$, out of the RS.
The $\Gamma_{j}$ are computed by starting the system in a single-photon Fock state in one of the RS QMs, which we then expand in the basis of the full-system QMs. We then fit calculated time evolution of the photon number in the chosen RS QM to an exponential. From this procedure, we find that the loaded decay rate of the pump RS QM, which was not altered with a coupling cavity, is $\Gamma_{P}\approx 5.30\times 10^{-5}$ $2\pi c/a$, and the intrinsic decay rate of the pump RS QM is $2\gamma_{P}\approx 1.63\times 10^{-5}$ $2\pi c/a$. The loaded decay rates of the signal and idler RS QMs are $\Gamma_{S}\approx 2.58\times 10^{-5}$ and $\Gamma_{I}\approx 2.6\times 10^{-5}$ $2\pi/c$ respectively, while their intrinsic decay rates are $2\gamma_{S}\approx 7.37\times 10^{-6}$ and $2\gamma_{I}\approx 2.38\times 10^{-5}$ $2\pi c/a$ respectively. The loaded decay rates of the signal and idler RS QMs were tuned to be approximately equivalent and close to half the loaded decay rate of the pump RS QM. Tuning was accomplished by introducing a coupling cavity (labeled in Fig. 1) between the RS and each output CROW. These coupling cavities have the effect of decreasing the coupling between the RS and their respective CROWs - in this way, we decreased the couplings independently until they were at the desired values.
Figure 7: Pump strength and numerical solutions to Eqs. (39-41) for a 6 pJ Gaussian pump pulse with a $k$-space width of $\kappa=2\pi/20d$. (a) Pump strength $g(\tilde{t})$ (solid green, left axis) and photon number in the signal and idler RS QMs (right axis). The upper curves are the total photon number in the signal (blue, solid) and idler (red, dashed) RS QMs. The lower curves are the thermal photon numbers in the signal (blue, solid) and idler (red, dotted) RS QMs. Note that the results are almost identical for the signal an idler modes. (b) Squeezing amplitude (orange, right axis) and correlation variance envelope (black, left axis) between the signal and idler RS QMs.
Eqs. (39-41) are solved numerically using a fourth-order Runge-Kutta method. In Fig. 7 (a), we plot the time evolution of the pump strength in the RS, $g(\tilde{t})$ (solid green), when a Gaussian pump pulse is initiated in the pump CROW with a pulse energy of 670 fJ, a center frequency at $\omega_{P}$, and $k$-space width parameter $\kappa=2\pi/20d$. For this pump, the peak number of pump photons in the RS is $\approx 2.3\times 10^{6}$. In Fig. 7 (a), we also show the calculated time evolution of the total photon number in the signal (red dashed) and idler (blue solid) RS QMs, and the calculated thermal photon number in the signal (red dotted) and idler (blue solid) RS QMs. We see that most of the generated photons are contributing to squeezing rather than to thermal noise. We note that the expectation value of the photon number in the generated entangled two-mode squeezed thermal state (2MSTS) is less than one, indicating the undepleted pump approximation is justified.
$+\expectationvalue{\quantity[\Delta\quantity(Y_{S}+Y_{I})]^{2}},$ (43)
where
are the quadrature operators for the $j^{\text{th}}$ RS QM. In these expressions, $\theta_{S}(t)$ and $\theta_{I}(t)$ are time dependent phases that can be chosen as desired. The correlation variance is an indicator of inseparability: if $\Delta^{2}_{S,I}<1$, then the state is entangled [38]. After substituting Eqs. (44, 45) into Eq. (III.3) we obtain
$\Delta^{2}_{S,I}$ $=\quantity(n_{S}^{\text{th}}+n_{I}^{\text{th}}+1)$ (46)
×$\times\quantity{\cosh(2r)+\cos(\theta_{S}+\theta_{I}+\theta)\sinh(2r)}.$
To isolate the minima (maxima) of the CV we set the argument of the second cosine to $\pi$ $(0)$ by selecting appropriate $\theta_{S}(t)$ and $\theta_{I}(t)$ such that $\theta_{S}+\theta_{I}+\theta=\pi\ (0)$, where $\theta$ is the phase of the squeezing parameter given in Eq. (38). We thus obtain expressions for the maximum and minimum CV in the between the signal and idler modes of the RS:
In Fig. 7 (b), we plot the time evolution of the squeezing amplitude, $r(\tilde{t})$ (orange), and the minium and maximum of the correlation variance in the RS (black). As can be seen, the CV reaches a minimum value of $\approx 0.2975$ at time $\tilde{t}_{0}\approx 0.9428$.
We now wish to calculate the correlation variance between a single-defect mode in the signal CROW and one in the idler CROW. To accomplish this, we select the time, denoted $\tilde{t}_{0}$ and labeled in Fig. 7, at which $\Delta_{S,I;\text{min}}^{2}$ is minimized. At this time, we retrieve the calculated quantities $r(\tilde{t}_{0})$, $n_{S}^{\text{th}}(\tilde{t}_{0})$, and $n_{I}^{\text{th}}(\tilde{t}_{0})$, given in Table 2, to define the density operator of the generated two-mode squeezed thermal state at the time $\tilde{t}_{0}$ according to Eq. (34). Note that, because we have designed the structure such that the signal and idler decay times are nearly equal and are much longer than the pump pulse duration in the RS, the number of thermal photons is very low and is much smaller than the total number of photons.
  $r(\tilde{t}_{0})$   $n_{S}^{\text{th}}(\tilde{t}_{0})$   $n_{I}^{\text{th}}(\tilde{t}_{0})$
0.6497 0.0458 0.0453
$\Delta_{S,I;\text{min}}^{2}(\tilde{t_{0}})$ $n_{S}(\tilde{t_{0}})$ $n_{I}(\tilde{t_{0}})$
0.2975 0.5749 0.5744
In the next section, we use this density operator as the starting point to compute the quantum evolution of the observables of interest in the signal and idler CROWs for times $\tilde{t}\geq\tilde{t}_{0}$. In doing this, we are assuming that the effect of the pump can be ignored for times $\tilde{t}>\tilde{t}_{0}$. This is a conservative approximation: Eqs. (39) and (42) show that the correlation variance will always be larger (worse) when no pump is present.
III.4 Correlation Variance in in the output CROWs

### v3

@@ Appendix A Quasi-Mode Bases
In this appendix we describe in detail the calculation of the various bases used in our theoretical approach. As our first step, we calculate the QMs of the individual defects and present the tight-binding formalism that we use to obtain the leaky QMs of the entire PC CCS discussed earlier and shown in Fig. 1. We first discuss some general properties of PC slabs and defect modes, and then show how we determine the QMs of a general system of coupled defects. We then compute the QMs of the full system and each of the subsystems.
A.0.1 General Procedure
@@
Defect-free PC slabs are periodic dielectric structures which, if properly designed, can support a photonic band gap - a set of frequencies for which light cannot propagate in the PC - similar to an electronic band gap in which no electronic states exist [PCBook]. Both of these band gaps arise from the periodicity of the underlying structure. The photonic band structure (and thus the size and location of the photonic band gap) of a PC slab is easily calculated using finite-difference time-domain (FDTD) simulations or finite-element analysis (FEA) by exploiting the high symmetry in a defect-free PC slab.
@@
One may form single-defect cavities in a PC slab by removing or shrinking an air hole at a particular lattice site, and thus locally disrupting the periodicity of the structure. Defects of this type are displayed in Fig. 1. These single-defect cavities in a square-lattice PC slab support a quadrupole-like resonant QM localized to the lattice site of the altered air hole [Fussell2007]. The resonant frequency of the single-defect cavity modes formed in this manner lies within the transverse electric photonic band gap of the underlying PC slab.
@@
The positive frequency part of the electric field of light in the $q^{\text{th}}$ individual defect QM is given by [KamandarDezfouli2014]
@@
$\curl\curl\textbf{M}_{q}(\textbf{r})-\frac{\tilde{\Omega}_{q}^{2}}{c^{2}}\epsilon_{q}(\textbf{r})\textbf{M}_{q}(\textbf{r})=0,$ (42)
@@
Figure 13: Single-defect QM frequency and quality factor as a function of defect air hole radius for the chosen PC. Computed with ANSYS Lumerical FDTD [FDTD].
The $\textbf{M}_{q}(\textbf{r})$ act as a basis for full system QMs that have frequencies well inside the chosen band gap of the PC. We use a tight-binding approach and expand the system QMs, $\textbf{N}_{\mu}(\textbf{r})$, in terms these as [Fussell2008]
@@
$\curl\curl\textbf{N}_{\mu}(\textbf{r})-\frac{\tilde{\omega}_{\mu}^{2}}{c^{2}}\epsilon(\textbf{r})\textbf{N}_{\mu}(\textbf{r})=0,$ (45)
@@
Eqs. (42), (43), and (45) lead to the generalized eigenvalue problem [Fussell2008]
@@
[…]
@@
where $V_{\mu}$ is the $\mu^{\text{th}}$ column of $\stackrel{{\scriptstyle\leftrightarrow}}{{V}}$. The $O_{\mu\nu}$ are elements of the square Hermitian matrix O, called the QM overlap matrix. We plot the overlaps between the full-system QMs in Fig. 5, and discuss their implications in Sec. III.
A.0.2 Resonant Structure Quasi-Modes
@@
$\curl\curl\textbf{C}_{j}(\textbf{r})-\frac{\tilde{\omega}_{j}^{2}}{c^{2}}\epsilon_{\text{RS}}(\textbf{r)}\textbf{C}_{j}(\textbf{r})=0,$ (53)
@@
In this case the RS is constructed from three single-defect cavities arranged in a line (see Fig. 1). Following Section A.0.1, the tight-binding eigenvalue problem for the RS QMs is
@@
A.0.3 Bloch Modes of an Infinite CROW
The remaining subsystems are the three CROWs. To understand the modes of the finite CROWS in the system, it is useful to first determine the Bloch QMs of infinite CROWs. In the nearest-neighbor tight-binding (NNTB) approximation, these modes can be expanded in the basis of single-defect cavity QMs as [Fussell2008]
@@ A.0.3 Bloch Modes of an Infinite CROW
where $d=3a$ is the CROW period and $N$ is the number of cavities in the CROW, which is taken to be of length $L=Nd$. In the limit that $N\rightarrow\infty$, the solution to the generalized eigenvalue problem for the CROW in the NNTB gives [Seifoory2019]
@@
Appendix B Pump Evolution
In this appendix, we describe how we model the evolution of the pump light in the full system and calculate the coherent state amplitude, $\alpha_{P}(t)$, of the pump in the RS pump mode. The pump at time $t=0$ starts in a coherent (classical) state in the pump CROW. We choose the initial state of the pump to be a Gaussian in the Bloch modes, centered at $k_{0}=-\frac{\pi}{2d}$. With this choice, the propagation speed is maximized, there is no waveguide group velocity dispersion, and the frequency is centered at the pump frequency of the RS: $\omega_{k_{0}}=\Omega_{0}=\omega_{P}$. Thus, the positive frequency part of the electric field of the pump light at time $t=0$ is given by
@@ Appendix B Pump Evolution
Now, to determine the time evolution of the pump in the full system, we need to expand the pump field at $t=0$ in terms of the QMs of the full system, since we know that these evolve in time according to Eq. (44). We start by inverting Eq. (43) to obtain the single cavity modes in terms of the full system modes:
@@
The time-dependent coherent state amplitude, $\alpha_{P}(t)$, in the RS pump mode is obtained by projecting $\textbf{E}_{P}^{+}(\textbf{r},t)$ onto the pump QM of the RS to obtain
@@
We note that the second line is approximate since in the previous line we use $\epsilon_{\text{RS}}$, while $\stackrel{{\scriptstyle\leftrightarrow}}{{B}}$ was computed using $\epsilon$.
Appendix C Nonlinear Generation in the Resonant Structure
In this appendix, we present the coupled differential equations used to determine the state parameters of the STS in the RS. This follows the work of Vendromin and Dignam [Vendromin2021]. Writing the nonlinear driving term as
@@ Appendix C Nonlinear Generation in the Resonant Structure
we find, for our choice of initial pulse in the structure we have designed, that $\beta$ is independent of $t$ over the time range for which the pump is non-negligible. When $\beta$ is independent of $t$ and the initial state is the vacuum state, it can be shown that the squeezing phase evolves simply according to [Vendromin2021]
@@
and the squeezing amplitude and thermal photon numbers are governed by the following set of coupled first-order differential equations [Vendromin2020]:
@@
where we employ the dimensionless time units $\tilde{t}=\Gamma_{+}t$ with $\Gamma{\pm}=(\Gamma_{S}\pm\Gamma_{I})/2$. The parameter $\zeta=\Gamma_{-}/\Gamma_{+}$, quantifies the difference in the loss rates in the two modes and is the pump strength defined in Eq. (15). Eqs. (69) to (71) are solved numerically using a fourth-order Runge-Kutta method, with the initial conditions: $r(0)=0$, $n^{\text{th}}_{S}(0)=0$, $n^{\text{th}}_{I}(0)=0$.
Appendix D Quantum State Evolution in the Output CROWs